User Tools

Site Tools


Telescope stuck in limits after the first change.. Had to be restarted and ok from about 10:23 UT.

Got drive error at 11:10:42, so observations will fail at this point. Caught the error and corrected at 11:40.

Got drive error at 16:03:43, caught the error and corrected at 16:22.

Another drive error at 16:29:18, corrected at 16:32.

Another drive error at 20:06:12, corrected at 20:08.

lbaops/lbafeb2014/v256daklog.txt · Last modified: 2015/12/18 16:38 by